Two Identical conducting spheres, fixed in place,
Answer #1

F = kQq/r²

0.109 = (8.99 * 10^9)Qq/0.713²

Qq = 6.163* 10^(-12)

But since we are told that the charges attract one another, we know that q1 and q2 have
opposite signs and so their product must be negative.

Qq = -6.163 * 10^(-12)

Q = -6.163 * 10^(-12)/q. . . . . . . . . . . . . . . (equation 1)


Then the two spheres are joined by a wire, If the new charge on each sphere is Q',

Q' + Q' = Q + q = 2Q'

F = kQ'q/r²

0.0441 = (8.99 * 10^9)(Q')²/0.713²

Q'=1.579 * 10^(-6)


Q + q = 2Q'

Q + q = 1.579 * 10^(-6). . . . . . . . . . . . . . . (equation 2)

x=_ 1.81555×10-6 and У= 3.39455×10-6 x= 3.39455×10-6 and y=-1.8 1555x10-6

both in C